OpenStudy (anonymous):

solve the equation log^(5)-(x+2)-log(5)6 =log(5)36

OpenStudy (anonymous):

\[\log_5 -(x+2)-\log_5 6=\log_5 36\]???is this ur question?

OpenStudy (anonymous):

i mean\[\log_5 (-(x+2))-\log_5 6=\log_5 36\]

OpenStudy (anonymous):

If so, raise 5 to both of them. \[5^{\log_5(-x-2)-\log_56}=5^{\log_5 36}\]

OpenStudy (anonymous):

\[\huge {5^{\log_5(-x-2)}*5^{-\log_5(6)}=5^{\log_5 36}}\] \[\huge {5^{\log_5(-x-2)}*(5^{\log_5(6)})^{-1}=5^{\log_5 36}}\] \[\large{\frac{-x-2}{6}=36}\]

OpenStudy (anonymous):

If not, use same method with actual question